A video camera is placed on a 120 m tall building. The angle of depression from the camera to the base of another building is 36o. The angle of elevation from the camera to the top of the same building is 47o.
How far apart are the two buildings? Round your answer to the nearest meter. Show work.

How tall is the building viewed by the camera? Round your answer to the nearest meter. Show work.

Answer by mananth(16946)


.
tan 36 = 120/AB
AB = 120/tan 36
AB =165 m
tan 47 = x/ 165
x= 165 * tan 47 =177
Height of building = 120+177= 297 m